# DexGarmentLab: Dexterous Garment Manipulation Environment with Generalizable Policy

### v2

@@ 2.1 Deformable Object Simulation
Simulator Scene Garment Sim on GPU Multiple Dexhands DexHand Garment Task Data Collection Method Demonstration Data Physically Plausible Robot-Garment Interaction
@@ 3.2 DexGarmentLab Physcial Simulation
Simulation Method. To achieve realistic simulation, we employ methods tailored to the physical properties of garments. Large garments (e.g., tops, dresses, trousers, etc.) are simulated using Position-Based Dynamics (PBD), while small, elastic items (e.g., gloves, hats) are modeled via the Finite Element Method (FEM). We provide detailed introduction and selection reason about PBD and FEM in Appendix F.1. Human avatars are represented by articulated skeletons with rotational joints and a skinned mesh for lifelike rendering.
@@ 3.4 DexGarmentLab Tasks
Dexterous (especially bimanual) garment manipulation is vital for domestic applications, yet it has not been thoroughly explored in existing research. To address this, we introduce 15 tasks across 8 garment categories (Fig. 2). Further details on these tasks are available in Appendix H.
@@ 4.2 Automated Data Collection Pipeline
Once target grasp points are identified on the operated garment using GAM, with demo task sequences, we control the robotic arms via inverse kinematics (IK) to execute sequential operations while controlling dexterous hands using PD controller based on the joint positions from demo hand grasp poses. It is important to note that the task trajectories are not fixed but are adapted based on the garment’s shape and length. For example, in garment-self-interaction tasks, such as folding, the lifting height is adjusted according to the sleeves and overall garment length to ensure proper folding. In garment-environment-interaction tasks, such as hanging, both the lifting height and placement position are adapted to align the garment’s center with the hanger, preventing slippage. These adjustments reflect common and reasonable actions in real-world scenarios, and introducing such variations increases the task difficulty. The details about tasks can be found in Appendix H.
During task execution, we can simultaneously record various information (such as images, point clouds, robot joint states, etc.) within the simulation environment. These serve as expert demonstration data for subsequent offline training of the policy. Details about recorded information can be found in Appendix C.
@@ 5 Generalizable Policy
Due to the poor generalization ability of current mainstream methods such as DP and DP3 for complex and variable garment manipulation scenarios, we propose Structure-Aware Diffusion Policy (SADP), a garment-environment-generalizable diffusion policy that ensures the generalization for different garment shapes and scene configurations, thereby enabling the smooth generation of subsequent trajectories after moving to the target manipulation area guided by GAM.
@@
Through experimental validation, SADP exhibits better generalization capabilities. We will further illustrate this advantage with experimental results in Sec. 6.2. Training details can be found in Appendix E.
@@ 6.1 Environment Setup
Tasks and Environments. We evaluate our method on 14 garment manipulation tasks with varying deformation characteristics. Detailed environment specifications including scene randomization parameters, success metrics, and train/test configurations are provided in the Appendix H.
Demonstration Collection. Using our automated data collection pipeline (Sec. 4), we acquired 100 demonstrations per task, with 30-100 environment steps in each demonstration. The time required to collect a demonstration varies between approximately 30-60 seconds depending on the task, which is significantly faster than data collection through teleoperation.
Baselines. We compare against two state-of-the-art diffusion-based approaches: 1) Diffusion Policy (DP)  [9], which utilizes images as observations to generate actions via diffusion. 2) 3D Diffusion Policy (DP3) [39], which replaces image inputs in the Diffusion Policy with point clouds.
@@ 6.2 Results Analysis
Tab. 2 (Row 1, 2, 5) quantifies HALO’s performance against baseline methods in simulation. Our method achieves superior success rates across dexterous garment manipulation tasks, demonstrating statistically significant improvements over existing approaches.
@@ 6.3 Real-World Experiments
Setup. Our setup comprises two RealMan RM75-6F (Arms) with Psibot G0-R (Dexterous Hands) and a RealSense D435 camera (As shown in Appendix D). We use Segment-Anything-2 [28] to segment the garment and interaction object from the scene and obtain corresponding point clouds.
@@
Figure 9: Real-World Experiments. We select Fold_Tops (Garment-Self-Interaction), Hang_Tops, Wear_Scarf, Wear_Hat (Garment-Environment-Interaction) as typical tasks for illustration. The test samples have different shapes, length, deformations while position of garments and interaction objects are variable. Despite this, GAM of HALO ensures accurate manipulation area location for grasp while SADP of HALO ensures adaptive trajectories base on garment and interaction object.
@@ 7 Conclusion
In this paper, we introduce DexGarmentLab, the first simulation environment designed to address the challenges of dexterous (especially bimanual) garment manipulation. Our work mainly makes three key contributions, including DexGarmentLab Environment, Automated Data Collection and Generalizable Policy. Through extensive experiments, we demonstrate that our approach can effectively learn complex manipulation tasks with minimal supervision and generalize across a wide range of garment shapes and deformation states in both simulation and real-world environments. The limitation of our work is discussed in Appendix F.
@@ B LeapMotion for Teleoperation
C Recorded Information in Automated Data Collection
@@
D Real-World Experiment Scene
@@
D Training Details of Main Algorithm
@@ G DexGarmentLab Assets
H Detailed Task Description
@@ Appendix B LeapMotion for Teleoperation
Here we just utilize the 24 finger joint positions to control Shadow Hand’s movement. The teleoperation system and visualization performance are shown in Fig 10. We provide relevant tutorial about how to use LeapMotion in our released code.
@@
Appendix C Recorded Information in Automated Data Collection
@@
Appendix D Real-World Experiment Scene
@@
Appendix E Training Details of Main Algorithm
E.1 Garment Affordance Model (GAM)
@@
E.2 Structure-Aware Diffusion Policy (SADP)
@@ Appendix F Limitation
F.1 Garment Simulation Method Limitation
@@
F.1.1 Position-Based-Dynamics (PBD)
@@
F.1.2 Finite-Element-Method (FEM)
In the simulator, garments modeled using Finite Element Method (FEM) are represented as a collection of discrete blocks, as shown in Figure  12. This approach effectively simulates non-deformable and elastic objects but does not accurately capture garment deformations. When garments undergo deformation in the simulator, they tend to revert to their original shape easily, a phenomenon that is particularly pronounced for larger garments. As shown in Fig. 12, when garments modeled using FEM are dragged, the discrete blocks used for simulation remain tightly connected and do not separate. Besides, the garment tends to preserve its original shape throughout the dragging process.
@@ F.1.2 Finite-Element-Method (FEM)
F.2 Environment Task Limitation
@@
F.3 Automated Data Collection Limitation
Compared to previous Automated Data Collection / Generation Research, we don’t compare our collected data with teleoperated data for further evaluation.
However, We want to clarify that we propose automated data collection to avoid the inefficiency and labor-intensive of teleoperation and supply data foundation for policy training. Beside, all the training data are collected in simulation environment rather than being generated artificially. And through the validation process and results we can also notice the trajectories output by trained policy is stable while the success rate of trained policy is reasonable, which implies the data collected by our automated data collection pipeline for policy training is reliable.
Appendix G DexGarmentLab Assets

### v3

@@ 2.1 Deformable Object Simulation
Simulator
Scene
Garment
Sim on
GPU
Multiple
Dexhands
DexHand
Garment Task
Data Collection
Method
Demonstration
Data
Physically Plausible
Robot-Garment Interaction
@@ 3.2 DexGarmentLab Physcial Simulation
Simulation Method. To achieve realistic simulation, we employ methods tailored to the physical properties of garments. Large garments (e.g., tops, dresses, trousers, etc.) are simulated using Position-Based Dynamics (PBD) [27], while small, elastic items (e.g., gloves, hats) are modeled via the Finite Element Method (FEM) [5]. We provide detailed introduction and selection reason about PBD and FEM in Appendix I.1. Human avatars are represented by articulated skeletons with rotational joints and a skinned mesh for lifelike rendering.
@@ 3.4 DexGarmentLab Tasks
Dexterous (especially bimanual) garment manipulation is vital for domestic applications, yet it has not been thoroughly explored in existing research. To address this, we introduce 15 tasks across 8 garment categories (Fig. 4). Further details on these tasks are available in Appendix K.
@@ 4.2 Automated Data Collection Pipeline
Once target grasp points are identified on the operated garment using GAM, with demo task sequences, we control the robotic arms via inverse kinematics (IK) to execute sequential operations while controlling dexterous hands using PD controller based on the joint positions from demo hand grasp poses. It is important to note that the task trajectories are not fixed but are adapted based on the garment’s shape and length. For example, in garment-self-interaction tasks, such as folding, the lifting height is adjusted according to the sleeves and overall garment length to ensure proper folding. In garment-environment-interaction tasks, such as hanging, both the lifting height and placement position are adapted to align the garment’s center with the hanger, preventing slippage. These adjustments reflect common and reasonable actions in real-world scenarios, and introducing such variations increases the task difficulty. The details about tasks can be found in Appendix K.
During task execution, we can simultaneously record various information (such as images, point clouds, robot joint states, etc.) within the simulation environment. These serve as expert demonstration data for subsequent offline training of the policy. Details about recorded information can be found in Appendix D.
@@ 5 Generalizable Policy
Due to the poor generalization ability of current mainstream methods such as DP and DP3 for complex and variable garment manipulation scenarios, we propose Structure-Aware Diffusion Policy (SADP), a garment-environment-generalizable diffusion policy that improves the generalization for different garment shapes and scene configurations, thereby enabling the smooth generation of subsequent trajectories after moving to the target manipulation area guided by GAM.
@@
Through experimental validation, SADP exhibits better generalization capabilities. We will further illustrate this advantage with experimental results in Sec. 6.2. Training details can be found in Appendix H.
@@ 6.1 Environment Setup
Tasks and Environments. We evaluate our method on 14 garment manipulation tasks with varying deformation characteristics. Detailed environment specifications including scene randomization parameters, success metrics, and train/test configurations are provided in the Appendix K.
Demonstration Collection. Using our automated data collection pipeline (Sec. 4), we acquired 100 demonstrations per task, with 30-100 environment steps in each demonstration. The time required to collect a demonstration varies between approximately 30-80 seconds depending on the task, which is significantly faster than data collection through teleoperation. We make comparison between autonomously collected data and teleoperation data in Appendix F.
Baselines. We compare against two state-of-the-art diffusion-based approaches across all the garment manipulation tasks: 1) Diffusion Policy (DP)  [10], which utilizes images as observations to generate actions via diffusion. 2) 3D Diffusion Policy (DP3) [41], which replaces image inputs in the Diffusion Policy with point clouds. What’s more, we have also additionally included four new baselines for comparison: ACT (IL), pi0 (VLA), RDT (VLA), and Eureka (RL+VLM) and select representative tasks for evaluation, including Fling Dress, Fold Trousers, Hang Coat, and Wear Bowlhat in simulation, as well as Fold Tops in real world, which can be found in Appendix G.
@@ 6.2 Results Analysis
Tab. 2 (Row 1, 2, 5) quantifies HALO’s performance against baseline methods in simulation. Our method achieves superior success rates across dexterous garment manipulation tasks, demonstrating statistically significant improvements over existing approaches. We also provide more baselines comparison in Appendix G.
@@ 6.3 Real-World Experiments
There are two ways using HALO for garment manipulation tasks:
1) transfer the data collection pipeline from simulation to real world, conduct automated data collection and policy training directly in the real world.
2) train the policy in the simulation and transfer the policy to the real world.
6.3.1 Experiments and Results Analysis for Way 1
we adopt way 1 to demonstrate the effectiveness of HALO. In the Appendix F Tab. 5, we report the efficiency and success rate of real-world data collection, highlighting the strong sim-to-real performance of GAM.
Figure 9: Real-World Experiments (Way 1). We select Fold_Tops (Garment-Self-Interaction), Hang_Tops, Wear_Scarf, Wear_Hat (Garment-Environment-Interaction) as typical tasks for illustration. The test samples have different shapes, length, deformations while position of garments and interaction objects are variable. Despite this, GAM of HALO ensures accurate manipulation area location for grasp while SADP of HALO ensures adaptive trajectories base on garment and interaction object.
Setup. Our setup comprises two RealMan RM75-6F (Arms) with Psibot G0-R (Dexterous Hands) and a RealSense D435 camera (As shown in Appendix E). We use Segment-Anything-2 [30] to segment the garment and interaction object from the scene and obtain corresponding point clouds.
@@ 6.3.1 Experiments and Results Analysis for Way 1
6.3.2 Experiments and Results Analysis for Way 2
Sim-to-real transfer from the simulation environment remains an important aspect of our study. To this end, as shown in Fig. 10, we aligned the settings of both the simulation and real-world environments by using the same hardware setup—Shadow Hand and UR10e—and selected two tasks, Hang Trousers and Wear Hat, for policy-level sim-to-real transfer, which means way 2. The evaluation criteria follow those used in the previous real-world experiments.
It is worth noting that sim-to-real performance is more sensitive to point cloud noise. To address the limited precision of the Realsense D435 in this context, we employed a Kinect camera for more accurate point cloud acquisition.
Figure 10: Real-World Experiments (Way 2). Two UR10e paired with ShadowHands and an Azure Kinect camera are used in our sim-to-real experiment. For each task, we show the scene configuration, affordance generated by GAM and trajectories generated by SADP.
Table 4: Performance Impact of Adding Real-World Data to Simulation Data
Task Name Only Simulation Data Simulation Data + 15 Real-World Data
Hang Trousers 8 / 15 (53.3%) 13 / 15 (86.7%)
Wear Hat 9 / 15 (60.0%) 13 / 15 (86.7%)
Experimental results in Tab. 4 show that due to the gap between simulation and the real world, training the policy solely on simulated data leads to a drop in sim-to-real performance. Incorporating a small amount of real-world data into the training process can effectively enhance the policy’s generalization ability.
@@ 7 Conclusion
In this paper, we introduce DexGarmentLab, the first simulation environment designed to address the challenges of dexterous (especially bimanual) garment manipulation. Our work mainly makes three key contributions, including DexGarmentLab Environment, Automated Data Collection and Generalizable Policy. Through extensive experiments, we demonstrate that our approach can effectively learn complex manipulation tasks with minimal supervision and generalize across a wide range of garment shapes and deformation states in both simulation and real-world environments. The limitation of our work is discussed in Appendix I.
Acknowledgment
This project was supported by National Natural Science Foundation of China (62376006) and National Youth Talent Support Program (8200800081).
@@ B LeapMotion for Teleoperation
C Advantages of Dexterous Hands over Parallel Grippers
Detailed explanation on advantages of dexterous hands over parallel grippers.
D Recorded Information in Automated Data Collection
@@
E Real-World Experiment Scene
@@
F Comparison with Teleoperation Data
Detailed comparison between autonomously collected data and teleoperation data to demonstrate the efficiency and high quality of automated data collection pipeline.
G Additional Baseline Comparison
Additional baseline comparison for demonstrating the excellent performance of HALO.
H Training Details of Main Algorithm
@@ I Limitation
K Detailed Task Description
@@ Appendix B LeapMotion for Teleoperation
Here we just utilize the 24 finger joint positions to control Shadow Hand’s movement. The teleoperation system and visualization performance are shown in Fig 11.
We have provided relevant tutorial about how to use LeapMotion in our released code.We also explain details on how LeapMotion generates grasping poses below.
For single specific task category (e.g., Hang Coat), the hand poses required to perform specific actions in specific regions, can generally be treated as consistent. Moreover, the deformable nature of garments allows them to adapt naturally to the dexterous hand’s pose during manipulation. Therefore, for each specific task, we first define a set of task-specific hand poses (the number depends on the task; for example, Hang Coat requires both a closed grasping pose for the collar and an open pose for release). These poses are generated via teleoperation using LeapMotion, with physical feasibility considered during generation. In this way, the pose definition process inherently includes a manual filtering step to select the most suitable poses.
These task-specific hand poses are then transferred directly into the data collection process for that task category. That is, during data collection across different garments within the same task, the same predefined hand poses are reused.
@@
Appendix C Advantages of Dexterous Hands over Parallel Grippers
Due to current restrictions in simulation technology, certain complex tasks cannot yet be effectively implemented in simulated environments. However, across the existing tasks, dexterous hands still demonstrate clear advantages over parallel grippers in several aspects.
1. Our experimental results, particularly those real-world results, reveal significant limitations in two-finger grippers when handling thin or flat regions of garments. In such cases, the gripper frequently fails to achieve a stable grasp, with garments either slipping during manipulation or not being grasped at all. Many existing demonstrations involve garment folding with two-finger grippers rely on grasping the garment’s boundary, but these strategies also often leads to repeated failed attempts. In contrast, our system employing a five-finger dexterous hand demonstrates clear advantages. Owing to the hand’s larger operation range, greater surface contact, and coordinated multi-finger control, it can stably grasp various regions of a garment—not just its edges—and maintain a firm hold throughout the manipulation process. This capability greatly enhances the flexibility and robustness of robotic garment handling, freeing the system from constraints on graspable areas or predefined strategies.
2. Furthermore, the diverse hand postures offered by the dexterous hand allow us to tailor the manipulation strategy to the specific requirements of each task.
  • to accomplish a fine-grained task such as putting on a glove, the hand can adopt a pinch posture to precisely grasp the glove edge.
  • For stable garment pick-and-place operations, an open/close posture can be used to maximize contact and grip.
  • When placing folded clothes, a cradle posture helps preserve the folded structure.
  • To smooth wrinkles, the hand can switch to a smoothening pose that maximizes surface coverage.
These example poses are shown in the bottom-left corner of Figure 1 (main paper).
The dexterous hand not only replicates the functionality of two-finger grippers but also enables finer and more robust manipulation of garments.
3. Although current simulation still faces significant challenges in accurately modeling interactions between deformable objects and dexterous hands—limiting its ability to simulate complex tasks—these limitations are expected to be gradually addressed with the continued advancement of computer graphics and robotics technologies. As a result, the advantages of dexterous hands will become even more prominent in complex tasks such as assistive dressing or knot tying, which require flexible coordination among multiple fingers. We will continue to extend tasks and optimize the simulation framework based on DexGarmentLab, aiming to support a wider range of more complex dexterous-hand garment interaction tasks.
Appendix D Recorded Information in Automated Data Collection
@@
Appendix E Real-World Experiment Scene
@@
Appendix F Comparison with Teleoperation Data
Table 5: Automated Data Collection Time and Success Rate Across All Tasks
Task Name Collection Single Collection Task Name Collection Single Collection
Success Rate Time Success Rate Time
Fling Dress 92.6% (100/108) 0 min 56 s Hang Trousers 99.0% (100/101) 0 min 58 s
Fling Tops 90.1% (100/111) 0 min 55 s Store Tops 98.0% (100/102) 1 min 05 s
Fling Trousers 90.1% (100/111) 0 min 48 s Wear Baseball Cap 85.5% (100/117) 0 min 55 s
Fold Dress 84.7% (100/118) 1 min 06 s Wear Scarf 83.3% (100/120) 1 min 42 s
Fold Tops 82.0% (100/122) 1 min 08 s Wear Bowl Hat 95.2% (100/105) 0 min 53 s
Fold Trousers 82.6% (100/121) 0 min 56 s Fold Tops (Real-World) 90.9% (50/55) 0 min 50 s
Hang Coats 92.6% (100/108) 0 min 41 s Hang Tops (Real-World) 96.2% (50/51) 0 min 38 s
Hang Dress 90.9% (100/110) 0 min 48 s Wear Scarf (Real-World) 83.3% (50/60) 1 min 15 s
Hang Tops 91.7% (100/109) 0 min 46 s Wear Hat (Real-World) 93.8% (50/53) 0 min 36 s
Tab. 5 presents the data collection time and success rate across all tasks, which validates the effectiveness and efficiency of the automated data collection pipeline. To further verify the quality of data acquired via automation, three representative tasks were selected for data collection, model training, and evaluation, using both teleoperation data and autonomously collected data. Specifically, teleoperation data in simulation was gathered using LeapMotion, while an exoskeleton device was employed for real-world teleoperation data collection. The performance of policies trained on teleoperation data was then compared with that of policies trained on autonomously collected data, with the results summarized in Tab. 6.
Table 6: Performance Comparison of HALO with Different Data Sources
Method (Data Source) Hang Tops (Simulation) Wear Bowlhat (Simulation) Fold Tops (Real-World)
HALO (Automated Collection Data) 0.92±0.04 0.72±0.04 13 / 15
HALO (Teleoperation Data) 0.88±0.03 0.70±0.06 13 / 15
Experimental results indicate that policies trained on teleoperated and autonomously collected data achieve comparable performance during evaluation. However, from the perspective of human and time cost, teleoperation-based data collection is significantly more labor-intensive and time-consuming than automated data collection. It is also important to note that our data is not synthetically generated, but collected through real executions in either simulation or the real world, which means the demonstrations are physically realistic. The entire data collection pipeline is enabled by model inference and one-shot demonstrations.
Appendix G Additional Baseline Comparison
We have additionally included four new baselines for comparison: ACT (IL), pi0 (VLA), RDT (VLA), and Eureka (RL+VLM). We selected representative tasks for evaluation, including Fling Dress, Fold Trousers, Hang Coat, and Wear Bowlhat in simulation, as well as Fold Tops in real world. The evaluation protocol remains consistent with the original paper:
  • For simulation tasks, each task is evaluated over 50 episodes using three different random seeds. We report the success rates as Mean ± Std across all trials.
  • For real-world task, we evaluate using 3 distinct garments per category, each tested under 5 different initial deformations. We report the success rates as successful_trials / all_trials.
Table 7: Performance Comparison of More Baseline Methods on Typical Tasks
Method Fling Dress Fold Trousers Hang Coat Wear Bowlhat Fold Tops
(Simulation) (Simulation) (Simulation) (Simulation) (Real World)
DP (IL) 0.59±0.05 0.47±0.04 0.52±0.04 0.41±0.05 9/15
DP3 (IL) 0.51±0.03 0.54±0.07 0.58±0.04 0.55±0.04 8/15
ACT (IL) 0.35±0.02 0.49±0.06 0.43±0.04 0.51±0.03 7/15
pi0 (VLA) 0.69±0.01 0.52±0.06 0.72±0.01 0.59±0.02 10/15
RDT (VLA) 0.60±0.02 0.58±0.02 0.62±0.02 0.48±0.01 9/15
Eureka (RL+VLM) / / 0.16±0.03 0.08±0.02 /
HALO (Ours) 0.82±0.06 0.77±0.02 0.90±0.01 0.72±0.04 13/15
Analysis For Imitation-Learning-Based Methods:
We collect 100 demonstrations for each simulation task and collect 50 demonstrations for each real-world task, which are used for policy training (DP/DP3/ACT). A notable limitation of their performance lies in the inability to accurately grasp the target region of the garment, as well as the lack of fine-grained control during placement based on the garment’s shape(/state).
Analysis for VLA Methods:
We combine all the demonstrations of simulation and real-world tasks to fine-tune pi0 and RDT model. In terms of final performance, pretrained models such as pi0 and RDT outperform from-scratch approaches like DP, DP3, and ACT. However, compared to HALO, VLA-based models still exhibit clear limitations in accurately perceiving garment shape and state, and in executing precise grasp-and-place actions, resulting in a noticeable performance gap relative to HALO.
Analysis for RL-Based Methods:
We selected Eureka as the representative RL-VLM-combined method and conducted preliminary experiments on several tasks in simulation. While we observed some initial successful cases on simpler tasks such as Hang Coat and Wear Bowlhat, the overall success rate remained low. Additionally, due to the lack of robust parallelization support in the current Isaac Sim environment and the limited effectiveness of reward functions in handling long-horizon tasks like Fling and Fold, this baseline did not yield valid results across all tasks. Nevertheless, we are actively developing a multi-parallel deformable object manipulation environment based on Isaac Lab, and RL-based approaches for deformable manipulation will be a key focus of our future work.
Appendix H Training Details of Main Algorithm
H.1 Garment Affordance Model (GAM)
@@
H.2 Structure-Aware Diffusion Policy (SADP)
@@ Appendix I Limitation
I.1 Garment Simulation Method Limitation
@@
I.1.1 Position-Based-Dynamics (PBD)
@@
I.1.2 Finite-Element-Method (FEM)
In the simulator, garments modeled using Finite Element Method (FEM) are represented as a collection of discrete blocks, as shown in Figure  13. This approach effectively simulates non-deformable and elastic objects but does not accurately capture garment deformations. When garments undergo deformation in the simulator, they tend to revert to their original shape easily, a phenomenon that is particularly pronounced for larger garments. As shown in Fig. 13, when garments modeled using FEM are dragged, the discrete blocks used for simulation remain tightly connected and do not separate. Besides, the garment tends to preserve its original shape throughout the dragging process.
@@ I.1.2 Finite-Element-Method (FEM)
I.2 Environment Task Limitation
@@
I.3 Generalizable Policy Limitation
Based on both simulation and real-world experimental results, we summarize the following points regarding policy limitation:
1. For garments with occluded or highly deformed regions, the Garment Affordance Model (GAM) does not always predict manipulation points accurately. This can lead to failed or suboptimal grasps by the dexterous hand, impacting the final performance. Nonetheless, GAM performs reliably on most garments. For challenging cases, an initial unfold action is recommended to expose key regions.
2. The policy struggles with garments that deviate from standard geometries, such as asymmetrical designs (e.g., single-sleeved tops) or heavily adorned costumes, which distort the point cloud and affect both GAM and SADP performance.
Appendix J DexGarmentLab Assets